OpenStudy (anonymous):

I Will Give A Medal! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1.Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

Slope intercept form is the form y = mx + b where m is the slope and b is the intercept. Just make your function look like that by solving for y. 3y = 1200 - 2x \[y=\frac{ -2 }{ 3 }x+400\] So the slope is: \[\frac{ -2 }{ 3 }\] and the intercept is 400.

OpenStudy (anonymous):

@snomasnayr could you help me with these also 2.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4.Graph the function using one of the following two options below. One the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area.